# Learning to Learn Transferable Generative Attack for Person Re-Identification

### v1

@@ Abstract
Deep learning-based person re-identification (re-id) models are widely employed in surveillance systems and inevitably inherit the vulnerability of deep networks to adversarial attacks. Existing attacks merely consider cross-dataset and cross-model transferability, ignoring the cross-test capability to perturb models trained in different domains. To powerfully examine the robustness of real-world re-id models, the Meta Transferable Generative Attack (MTGA) method is proposed, which adopts meta-learning optimization to promote the generative attacker producing highly transferable adversarial examples by learning comprehensively simulated transfer-based cross-model&dataset&test black-box meta attack tasks. Specifically, cross-model&dataset black-box attack tasks are first mimicked by selecting different re-id models and datasets for meta-train and meta-test attack processes. As different models may focus on different feature regions, the Perturbation Random Erasing module is further devised to prevent the attacker from learning to only corrupt model-specific features. To boost the attacker learning to possess cross-test transferability, the Normalization Mix strategy is introduced to imitate diverse feature embedding spaces by mixing multi-domain statistics of target models. Extensive experiments show the superiority of MTGA, especially in cross-model&dataset and cross-model&dataset&test attacks, our MTGA outperforms the SOTA methods by 21.5% and 11.3% on mean mAP drop rate, respectively. The code of MTGA will be released after the paper is accepted.
@@
Person re-identification aims at retrieving specific persons from security surveillance systems [1, 2]. Along with the advancement of deep neural networks[3, 4, 5], it has made remarkable progresses and been widely applied to intelligent surveillance systems [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, it has been found that deep neural networks are vulnerable to adversarial attacks[16, 17, 18, 19, 20], which can mislead deep neural network models by adding imperceptible perturbations to benign images. Deep learning-based re-id models inevitably inherit the vulnerability of deep networks to adversarial samples[21, 22], which makes public safety under great threat. To study the security of surveillance systems, it is important to explore the vulnerability of the deep learning-based re-id models to adversarial samples.
@@ II-A Transferable Adversarial Attack
Szegedy et al.[16] found the intriguing transferability of attack examples, which permits attackers to generate adversarial example from surrogate models to attack black-box target models. Since this property of adversarial examples poses real-world DNN applications under serious security concerns, there have been extensive works aiming to improve the transferability of adversarial examples, which can be categorized into four groups, namely input transformation attacks [35, 36, 37, 38], gradient modification attacks [39, 40, 41], intermediate feature attacks [42, 43, 44] and model ensemble attacks [45, 46, 47]. Nevertheless, these approaches merely consider the cross-model transferability, presuming that the distribution of the attacked images and the target model training data are consistent, which can not been guaranteed in real-world situations. There are very few literatures focusing on this issue. Naseer et al. [48] trained a generative network that produces transferable cross-dataset perturbations by maximizing the fooling gap using relativistic supervisory signal. Zhang et al. [31] trained the attacker to disrupt low-level features and enhanced the transferability towards black-box domains by randomly normalizing benign images at image-level. Li et al. [30] trained a domain-agnostic feature extractor by self-supervised learning as the surrogate model to accomplish the cross-dataset transferability. Our MTGA is proposed for more complicated cross-model&dataset&test transferable attacks on black-box re-id models.
II-B Adversarial Attack Against Person Re-id
Previous attack methods [49, 50, 51] have mainly concentrated on the image classification task, aiming to significantly alter class predictions. But these methods are inapplicable to attack re-id models[23], since the re-id task is an open-set task. To effectively attack re-id models, Bai et al. [21], Zheng et al. [23] and Bouniot et al. [24] proposed different metric attack methods based on feature similarity calculation. Nevertheless, these works conducted white-box attacks, which need to know the parameters of target re-id models and are only valid for seen adversarial attack training data. They are not practical, because attackers need to attack unknown models and unseen queries in the realistic scenarios. To accomplish black-box attacks against re-id models, some researches [25, 27, 26, 28, 29] studied the transferable attacks for re-id system. Yang et al. [26] and Subramanyam [29] only enhanced the cross-dataset transferability by adopting multi-source datasets in additive and generative attack, respectively. Wang et al. [28] also ignored the cross-model transfer capabilities and developed a multi-stage discriminator network for cross-dataset general attack learning. Ding et al. [27] merely focused on cross-model transferability, introducing a model-insensitive regularization term for universal attack against different CNN structures. Yang et al. [25] built a combinatorial attack that consists of a functional color attack and universal additive attack to promote the cross-model&dataset of the attack. However, they attacked domain-specific target models, neglecting the diversity of feature embedding spaces of the same model. These existing transfer-based re-id black-box attacks have not yet achieved high transferability, without take into account the cross-test transferability of re-id adversarial examples.
@@ II-C Meta-learning
Meta-learning is a learning-to-learn[52] algorithm, which aims to improve further learning performance by distilling the experience from multiple learning episodes (i.e. meta-train and meta-test processes)[53, 54]. It has been widely used in deep learning tasks, e.g. few-shot learning, domain generalization and hyperparameter optimization. Recently, some meta-learning based transferable adversarial attack methods have been proposed, the underlying concept of which is to construct numerous meta transfer attack tasks. Yuan et al. [55] enhanced the cross-model transferability by composing different cross-model meta attack tasks. Fang et al. [56] composed transfer attack tasks with data augmentation and model augmentation, through randomized data transformation and model backpropagation altering. Yin et al. [57] generalized the generic prior of examples by treating attack on each examples as one task and fine-tuning the surrogate model during the meta-test process.
Distinct from above adversarial attack methods for re-id and meta-learning based attack methods, our method constructs extensive cross-model&dataset&test black-box adversarial attack tasks for attackers to learn how to generate more generic and transferable adversarial examples. And our CAS, PRE and NorMix modules are quite distinct from others.
@@ III-C Meta Task Generation
Cross-model&dataset Attack Simulation method. Because of the unknown parameters of the re-id model and unseen domain queries to be attacked in black-box scenarios, the adversarial generator needs to learn to handle the cross-model and cross-dataset attack situations. To mimic this case, Cross-model&dataset Attack Simulation method is proposed, which makes the target model and input data different during meta-train and meta-test process. Concretely, the data zoo and the model zoo that contains multiple datasets and multiple re-id models are constructed, from which CAS randomly selects distinct models and data for meta-train and meta-test processes to simulate cross-model and cross-dataset attacks. To represent numerous models well, CAS takes baseline models of three mainstream approaches (i.e. global-based, part-based and attention-based) to construct the model zoo.
@@
Perturbation Random Erasing strategy. Although there are several surrogate models in the model zoo to allow the attacker learning to handle cross-model attack scenarios, the number of these models is still limited, which may result in the attacker only learning to attack model-specific features. To address this problem, the Perturbation Random Erasing strategy is proposed to boost the attacker to disrupt holistic person features. PRE randomly selects a rectangle region of generated perturbations and erases it to generate incomplete perturbations. These incomplete perturbations prompt the attacker not to rely only on corrupting specific region features, as perturbations in these specific regions may be erased, leading to the failure of damaging specific region features. PRE is adopted in the meta-test process to test the attack error of trained adversarial attackers with generated incomplete perturbations, optimizing that error will enhance the attacker to achieve holistic destruction of image features and improve the transferability against black-box models.
@@
Normalization Mix module. The models that trained with different domain data could project person images to various feature embeddings, even though they share the same model architecture. NorMix is devised to project features to different feature embedding spaces, which is applied in meta-test process to promote the attacker learning to handle this cross-test issue. Specifically, there are multiple batch-norm layers [58] across the re-id model architectures, the statistics of which imply the distribution of the model training data. The batch normalization is formulated as
where is the input feature, and are the mean and variance of , and are learnable affine parameters used for linear transformation. To get diverse feature embeddings that the test data may be projected by the target model, the statistic of each batch-norm layer is mixed by
@@
where and are the empirical mean and variance of the pretrained meta-test model , and are the training statistics of the and is the mix coefficient that sampled from Beta Distribution. With mixed mean and variance , meta-test data features can be embedded to different feature spaces by
where and are copied from the batch-norm layers of meta-test model.
@@ III-D Optimization Procedure
The parameters of adversarial generator are supposed to be optimized by the meta-learning optimization. To disrupt the retrieval list of generated adversarial examples, the attacked image features should be far away from the original features. In our MTGA, the adversarial Euclidean Distance () loss
is applied to corrupt the similarity of adversarial features and benign features. Meanwhile, and are trained by the GAN loss respectively, denote as:
@@
IV Experiments
IV-A Experimental Setup
Model zoo and data zoo. Model zoo is composed of IDE[2], PCB[59] and ViT[13], which are all trained on the DukeMTMC[60] datasets. And the data zoo consists of DukeMTMC[60], CUHK03[61], and MSMT17[62] datasets.
Black-box re-id models and unseen queries. To evaluate the transferability of our adversarial generator to different re-id models, numerous re-id models (i.e. BOT[63], LSRO[64], MuDeep[65], Aligned[66], MGN[67], HACNN[68], Transreid[13], PAT[69]) are taken to act as the black-box re-id models. Notably, these models are in different backbones, including ResNet[3] (i.e. BOT[63]), ViT[4] (i.e. Transreid[13], PAT[69]), DenseNet[70] (i.e. LSRO[64]) and Inception-v3[71] (i.e. MuDeep[65]). Also, these models are in different architectures, including global-based (i.e. BOT[63]), part-based (i.e. MGN[67]) and attention-based (i.e. HACNN[68]). In order to test the transferabilities on different domain models, these models are trained on different domain datasets (i.e. Market[72] and DukeMTMC[60]). Meanwhile, to test the transferability of our attacker to unseen queries, VIPeR[73] and Market[72] datasets play the role of unseen domain data.
@@ IV-A Experimental Setup
where is the aAP of the re-id models on the benign images and is on the generated adversarial examples.
Implementation Details. MAML[54] is adopted as our meta-learning framework and in each iteration meta-tasks are generated. Adam[74] optimizer is employed to optimize the model parameters. The learning rate of inner loop and outer loop are set to 1e-4 and 2e-4. The generator and discriminator model are referenced to the Mis-Ranking[28]. All experiments are performed by -bounded attacks with , where is the upper bound for the change of each pixel. The mix coefficient of NorMix is sampled from Beta Distribution, i.e., .
IV-B Comparison with State-of-the-art Methods
We compare our proposed MTGA method with state-of-the-art attack methods on transferable black-box re-id attacks, including MUAP[27], Mis-Ranking[28], MetaAttack[25]. These methods are all re-trained by attacking IDE[2] on DukeMTMC[60]. Unlike other methods, MetaAttack[25] method incorporates the color attack in addition to the additive perturbation. For a fair comparison, we only compare the attack performances of its additive perturbation. Meanwhile, based on these original methods, we train MetaAttack*, Mis-Ranking*, MUAP* and MTGA* in the ensemble training setting by attacking models in the model zoo (i.e. IDE[2], PCB[59] and ViT[13]) with dataset in data zoo (i.e. DukeMTMC[60], CUHK03[61], and MSMT17[62]). The experiment details of training data, surrogate model, target data and target model are shown in Tab. I. The comparison results on the mAP, aAP and mDR of six black-box attack settings are shown in Tab. III to Tab. VII.
Comparisons with original SOTA methods. It can be seen that in every black-box attack scenario, our MTGA performs much better than other SOTA methods on attacking multiple black-box re-id models. For most practical and challenging cross-model&dataset&test scenario, our MTGA achieves a superior performance of 18.5% aAP and 51.3% mDR score, which outperforms the SOTA methods by 4.3% and 11.3% in terms of aAP and mDR. For cross-model&dataset attack setting, our MTGA also gets the best transferability results, surpassing others by 16.5% and 21.5% in terms of aAP and mDR.

### v2

@@ Abstract
Deep learning-based person re-identification (re-id) models are widely employed in surveillance systems and inevitably inherit the vulnerability of deep networks to adversarial attacks. Existing attacks merely consider cross-dataset and cross-model transferability, ignoring the cross-test capability to perturb models trained in different domains. To powerfully examine the robustness of real-world re-id models, the Meta Transferable Generative Attack (MTGA) method is proposed, which adopts meta-learning optimization to promote the generative attacker producing highly transferable adversarial examples by learning comprehensively simulated transfer-based cross-model&dataset&test black-box meta attack tasks. Specifically, cross-model&dataset black-box attack tasks are first mimicked by selecting different re-id models and datasets for meta-train and meta-test attack processes. As different models may focus on different feature regions, the Perturbation Random Erasing module is further devised to prevent the attacker from learning to only corrupt model-specific features. To boost the attacker learning to possess cross-test transferability, the Normalization Mix strategy is introduced to imitate diverse feature embedding spaces by mixing multi-domain statistics of target models. Extensive experiments show the superiority of MTGA, especially in cross-model&dataset and cross-model&dataset&test attacks, our MTGA outperforms the SOTA methods by 20.0% and 11.3% on mean mAP drop rate, respectively. The source codes are available at https://github.com/yuanbianGit/MTGA.
@@
Person re-identification aims at retrieving specific persons from security surveillance systems [1, 2]. Along with the advancement of deep learning, it has made remarkable progresses and been widely applied to intelligent surveillance systems [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. However, it has been found that deep neural networks are vulnerable to adversarial attacks [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], which can mislead deep neural network models by adding imperceptible perturbations to benign images. Deep learning-based re-id models inevitably inherit the vulnerability of deep networks to adversarial samples [24, 25], which makes public safety under great threat. To study the security of surveillance systems, it is important to explore the vulnerability of the deep learning-based re-id models to adversarial samples.
@@ II-A Transferable Adversarial Attack
Szegedy et al. [14] demonstrated the transferability of adversarial examples, enabling attackers to craft examples on surrogate models to attack target black-box models. Efforts to enhance adversarial transferability can be grouped into four categories: input transformation [38], gradient modification [39], intermediate feature manipulation [40], and model ensemble strategies [41]. However, these methods focus solely on cross-model transferability, assuming consistent data distributions between attacked images and target model training data, which is rarely met in practical scenarios. Cross-dataset transferability has received limited attention. Naseer et al. [42] proposed a generative network to produce cross-dataset perturbations by maximizing the fooling gap. Zhang et al. [34] disrupted low-level features and improved transferability by randomly normalizing benign images. Li et al. [33] employed self-supervised learning to train a domain-agnostic feature extractor for cross-dataset attacks. Yang et al. [43] leveraged vision-language models and prompt learning to enhance cross-dataset transferability. In contrast, our MTGA is designed for more complex cross-model&dataset&test attacks targeting black-box re-id models.
II-B Adversarial Attack Against Open-set Task
Person re-id is a specialized image retrieval task focused on identifying a target individual across non-overlapping camera views [44, 45, 46, 47, 48, 49, 50]. Unlike classification tasks, re-id operates as an open-set problem, where the test classes differ entirely from the training classes [51]. Previous attack methods [52, 53, 54] on the image classification task are inapplicable to attack open-set task models [26]. To effectively attack open-set re-id, face recognition and image retrieval models, some white-box attack methods based on feature similarity [24, 26, 27] and rank results disruption [55, 56] have been developed. To accomplish black-box attacks against these models, researchers studied the transferable attacks on open-set tasks. Gong et al. [57] used more obvious color variation to randomly disturb the retrieved images. Yang et al. [29] and Subramanyam [32] enhanced the cross-dataset transferability by adopting multi-source datasets in additive and generative attack, respectively. Wang et al. [31] developed a multi-stage discriminator network for cross-dataset general attack learning. Ding et al. [30] introduced a model-insensitive regularization term for universal attack against different CNN structures. Yang et al. [28] built a combinatorial attack that consists of a functional color attack and universal additive attack to promote the cross-model&dataset of the attack. Zhong et al. [58] applied dropout layers to boost cross-model transferability. Li et al. [59] leveraged a highly related task as the sibling task to generate cross-model&dataset transferable attacks. Existing transfer-based open-set adversarial attack methods have incorporated considerations for diverse test data domains in open-set scenarios, along with cross-model architectural transferability. However, these approaches still fail to account for potential variations in the training domains of target models, which is particularly crucial for ensuring effectiveness in cross-test attack scenarios.
@@ II-C Meta-learning
Meta-learning is a learning-to-learn [60] algorithm, which aims to improve further learning performance by distilling the experience from multiple learning episodes (i.e., meta-train and meta-test processes) [61, 62]. It has been widely used in deep learning tasks, e.g., few-shot learning, domain generalization and hyperparameter optimization. Recently, some meta-learning based transferable adversarial attack methods have been proposed and show superiority to traditional attack method. Unlike traditional methods that train attacks on a single model or input, the meta-learning based method construct numerous meta transfer attack tasks using multiple models or inputs for training. They construct transferability error tests by meat-train and meta-test training tasks, instead of performing accuracy degradation attacks as in traditional methods, which makes the meta-learning based attack approaches to get better transferability. Yuan et al. [63] enhanced the cross-model transferability by composing different cross-model meta attack tasks. Fang et al. [64] composed transfer attack tasks with data augmentation and model augmentation, through randomized data transformation and model backpropagation altering. Yin et al. [65] generalized the generic prior of examples by treating attack on each examples as one task and fine-tuning the surrogate model during the meta-test process.
Distinct from above adversarial attack methods for open-set and meta-learning based attack methods, our method constructs extensive cross-model&dataset&test black-box adversarial attack tasks for attackers to learn how to generate more generic and transferable adversarial examples. And our CAS, PRE and NorMix modules are quite distinct from others.
@@ III-C Meta Task Generation
Cross-model&dataset Attack Simulation method. Because of the unknown parameters of the re-id model and unseen domain queries to be attacked in black-box scenarios, the adversarial generator needs to learn to handle the cross-model and cross-dataset attack situations. To mimic this case, Cross-model&dataset Attack Simulation method is proposed, which makes the target model and input data different during meta-train and meta-test process. Concretely, the data zoo and the model zoo that contains multiple datasets and multiple re-id models are constructed, from which CAS randomly selects distinct models and data for meta-train and meta-test processes to simulate cross-model and cross-dataset attacks. To represent numerous models well, CAS takes baseline models of three mainstream approaches (i.e., global-based, part-based and attention-based) to construct the model zoo.
@@
Perturbation Random Erasing strategy. Although there are several surrogate models in the model zoo to allow the attacker learning to handle cross-model attack scenarios, the number of these models is still limited, which may result in the attacker only learning to attack model-specific features. To address this problem, the Perturbation Random Erasing strategy is proposed. Base on the observation that different models tend to focus on distinct discriminative regions during recognition [38], our PRE aims to prevent attacker from perturbing only model-specific feature regions by adding randomly erased incomplete perturbations on images, thereby boosting the attacker to disrupt holistic image features and enhancing the cross-model transferability of adversarial attacks. Specifically, our PRE can be formulated by
where is the random pattern and is the Hadamard product. The random pattern is initially configured as a matrix of size with all elements set to 1, where and is the height and width of the person image . Then, with a probability for undergoing PRE, a randomly selected percentage of its elements are set to 0. Finally, the matrix is rescaled to the dimensions using nearest-neighbor interpolation to get the patch masked random pattern . By applying Hadamard product between the final random pattern and the perturbation, we can erase random patch regions of the perturbation. These incomplete perturbations prompt the attacker not to rely only on corrupting specific region features, as perturbations in these specific regions may be erased, leading to the failure of damaging specific region features.
With the PRE mechanisms, the gradient backpropagation for updating the adversarial generator parameters with Eq.8 loss can be formulated by:
It can be observed that our PRE, by introducing , enables the stochastic sparsification of gradients, thereby suppressing the learning on model-specific salient features and facilitating the disruption of comprehensive features of the agent model . From another perspective, the perturbations after random erasure can be regarded as perturbations generated by different generators. This is equivalent to training multiple generators, ultimately resulting in an implicit ensemble-averaged output, which reduces the variance of the attack effectiveness and renders the generated perturbations more generalizable.
PRE is adopted in the meta-test process to test the attack error of trained adversarial attackers with generated incomplete perturbations, optimizing that error will enhance the attacker to achieve holistic destruction of image features and improve the transferability against black-box models.
@@
Normalization Mix module. The models that trained with different domain data could project person images to various feature embeddings, even though they share the same model architecture. NorMix is devised to project features to different feature embedding spaces, which is applied in meta-test process to promote the attacker learning to handle this cross-test issue. NorMix is motivated by the insight that the weight matrix typically captures label information, whereas the Batch Normalization (BN) layer [66] houses domain-specific knowledge [67]. Building on this, we introduce the Normalization Mix module to simulate various feature embeddings by blending the BN statistics, which reflecting the underlying distribution of the model’s training data.
Specifically, there are multiple batch-norm layers across the re-id model architectures, and the batch normalization is formulated as
where is the input feature, and are the mean and variance of , β^f
@@
where and are the empirical mean and variance of the pretrained meta-test model , and are the training statistics of the meta-train datasets on and is the mix coefficient that sampled from Beta Distribution. Because the training dataset of the and the meta-train dataset are different, we can get effective mixed BN statistics by mixing them. With the dynamical mix coefficient and different meta-train datasets , diverse mixed mean and variance can be obtained. Finally, meta-test data features can be embedded to different feature spaces by
where and are copied from the batch-norm layers of meta-test model, is the output features mapped to diverse embeddings. By leveraging derived from various feature spaces throughout meta-test process, our MTGA can facilitate the adversarial generators learning cross-test transfer capability.
NorMix can also be viewed as a generalization of test-time adaptation (TTA) methods [68, 69], which aim to adapt models to unseen test domains in real-time by dynamically adjusting normalization statistics while preserving the original network parameters. TTA approaches address test and training data distribution shift by replacing the original BN parameters with statistics estimated from test batches. NorMix aligns with TTA’s principle that domain-specific knowledge resides primarily in BN layers, adapting models’ feature embeddings to diverse distributions by dynamically adjusting BN statistics while preserving network weights, thereby validating that normalization-layer adjustments can simulate cross-test feature distributions.
@@ III-D Optimization Procedure
The parameters of adversarial generator are supposed to be optimized by the meta-learning optimization. To disrupt the retrieval list of generated adversarial examples, the attacked image features should be far away from the original features. In our MTGA, the adversarial Euclidean Distance loss
is applied to corrupt the similarity of adversarial features and benign features extracted by the re-id model , where is the Euclidean distance. Meanwhile, and are trained by the GAN loss respectively, denote as:
@@
IV Experiments
To evaluate the superiority of our method, we first provide training and evaluation settings in the Section IV-A and then present experimental results in Section IV-B. Afterwards, comprehensive evaluations including ablation studies, adversarial example quality assessment, visualization analyses, and attack effectiveness against defense mechanisms are provided to further validate the efficacy of our proposed method.
IV-A Experimental Setup
Training details. Model zoo is composed of IDE [2], PCB [70] and ViT [12], which are all trained on the DukeMTMC [71] datasets. And the data zoo consists of DukeMTMC [71], CUHK03 [72], and MSMT17 [73] datasets. MAML [62] is adopted as our meta-learning framework and in each iteration meta-tasks are generated. Adam [74] optimizer is employed to optimize the model parameters. The learning rate of inner loop and outer loop are set to 1e-4 and 2e-4. The generator and discriminator model are referenced to the Mis-Ranking [31]. All experiments are performed by -bounded attacks with , where is the upper bound for the change of each pixel. The mix coefficient of NorMix is sampled from Beta Distribution, i.e., . The probability of undergoing PRE is set to and the mask percentage of random pattern is set to .
Evaluation models and datasets. To evaluate the transferability of our adversarial generator to different re-id models, numerous re-id models (i.e., BOT [75], LSRO [76], MuDeep [77], Aligned [78], MGN [79], HACNN [80], Transreid [12], PAT [81]) are taken to act as the black-box re-id models. Notably, these models are in different backbones, including ResNet [82] (i.e., BOT [75]), ViT [83] (i.e., Transreid [12], PAT [81]), DenseNet [84] (i.e., LSRO [76]) and Inception-v3 [85] (i.e., MuDeep [77]). Also, these models are in different architectures, including global-based (i.e., BOT [75]), part-based (i.e., MGN [79]) and attention-based (i.e., HACNN [80]). In order to test the transferabilities on different domain models, these models are trained on different domain datasets (i.e., Market [86] and DukeMTMC [71]). Meanwhile, to test the transferability of our attacker to unseen queries, VIPeR [87] and Market [86] datasets play the role of unseen domain data.
@@ IV-A Experimental Setup
where is the aAP of the re-id models on the benign images and is on the generated adversarial examples. Smaller aAP and larger mDR represent better transferability of adversarial examples, so we use aAP and mDR to indicate this relationship in the table of experimental results for clearer comparisons.
@@
IV-B Experimental Results
We compare our proposed MTGA method with state-of-the-art attack methods on transferable black-box re-id attacks, including MUAP [30], Mis-Ranking [31], MetaAttack [28], and also with state-of-the-art transferable generative attack methods, including GAP [88], CDA [42], LTP [89], BIA [90] and PDCL-Attack [43]. These methods are all re-trained by attacking IDE [2] on DukeMTMC [71]. Unlike other methods, MetaAttack [28] method incorporates the color attack in addition to the additive perturbation. For a fair comparison, we only compare the attack performances of its additive perturbation. Meanwhile, based on these original methods, we train MetaAttack*, Mis-Ranking*, MUAP* and MTGA* in the ensemble training setting by attacking models in the model zoo (i.e., IDE [2], PCB [70] and ViT [12]) with dataset in data zoo (i.e.,DukeMTMC [71], CUHK03 [72], and MSMT17 [73]). The experiment details of training data, surrogate model, target data and target model are shown in Tab. I. The comparison results on the mAP, aAP and mDR of six black-box attack settings are shown in Tab. III to Tab. VII.
Comparisons with original SOTA methods. It can be seen that in every black-box attack scenario, our MTGA performs much better than other SOTA methods on attacking multiple black-box re-id models. For most practical and challenging cross-model&dataset&test scenario, our MTGA achieves a superior performance of 18.5% aAP and 51.3% mDR score, which outperforms the SOTA methods by 4.3% and 11.3% in terms of aAP and mDR. For cross-model&dataset attack setting, our MTGA also gets the best transferability results, surpassing others by 15.4% and 20.0% in terms of aAP and mDR. It is noteworthy that generative attack methods, including GAP, CDA, LTP, and BIA, achieve moderate performance, likely because their classification loss functions are ill-suited for the re-id retrieval task. However, PDCL-Attack attains suboptimal adversarial transferability by leveraging a vision-language model to guide semantic disruption in images. Meta-Attack demonstrates superior cross-dataset transferability due to its incorporation of diverse datasets during training, though its performance remains limited in other scenarios. Methods such as Mis-Ranking and MUAP improve transferability by introducing multi-stage discriminator networks and model-insensitive regularization terms, achieving reasonable results. Nonetheless, these two methods neither explicitly optimize transferability as a primary objective nor comprehensively consider cross-test scenarios, resulting in performance that still falls short of our approach.